# Optimal Spectral Transitions in High-Dimensional Multi-Index Models

### v1

@@ C.1 State Evolution
${\mathcal{F}}({\bm{M}};a,{\bm{V}})$ $\coloneqq\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T}(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T})^{-1}{\bm{M}}(\mathrm{Cov}[\bm{\mathbf{z}}|y]-{\bm{I}})\right],$ (96)
@@
${\mathcal{G}}({\bm{M}};a,{\bm{V}})$ $\coloneqq\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T}(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T})^{-1}{\bm{M}}(\mathrm{Cov}[\bm{\mathbf{z}}|y]-{\bm{I}}){\bm{M}}^{T}(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T})^{-1}{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T}\right].$ (98)
@@
$\langle{\mathcal{F}}({\bm{M}};a,{\bm{V}}),{\bm{M}}^{\prime}\rangle$ $=\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}{\rm Tr}\left[(\mathrm{Cov}[\bm{\mathbf{z}}|y]-{\bm{I}}){\bm{M}}^{T}(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T})^{-1}{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T}{\bm{M}}^{\prime}\right]$ (99)
$=\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}{\rm Tr}\left[(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T})^{-1}{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T}{\bm{M}}^{\prime}(\mathrm{Cov}[\bm{\mathbf{z}}|y]-{\bm{I}}){\bm{M}}^{T}\right]$ (100)
$={\rm Tr}\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T}(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T})^{-1}{\bm{M}}^{\prime}(\mathrm{Cov}[\bm{\mathbf{z}}|y]-{\bm{I}}){\bm{M}}^{T}\right]$ (101)
@@
Furthermore, in order to bound the operator norm of ${\bm{V}}$, we choose $\gamma_{t}$ such that $\|{\bm{V}}_{t+1}\|_{\mathrm{op}}=1$.
@@
  • Informed fixed points. Choosing $a_{t}$ such that
@@
    where we used, introducing the auxiliary notation ${\mathcal{L}}(y,a,{\bm{V}})=(\mathrm{Cov}[\bm{\mathbf{z}}|y](a{\bm{V}}^{-1}-{\bm{V}})+{\bm{V}})^{-1}$
@@
    $=\mathbb{E}_{{\mathrm{y}}}{\rm Tr}\left[{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}})^{-1}{\bm{Q}}{\bm{V}}(\mathrm{Cov}[\bm{\mathbf{z}}|y]-{\bm{I}}){\mathcal{L}}({\mathrm{y}},a,{\bm{V}})^{T}\right]$
    $\leq\mathbb{E}_{{\mathrm{y}}}{\rm Tr}\left[{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}})^{-1}{\bm{Q}}{\bm{V}}(\mathrm{Cov}[\bm{\mathbf{z}}|y]-{\bm{I}})\right]\underbrace{{\rm Tr}\left[{\mathcal{L}}({\mathrm{y}},a,{\bm{V}})^{T}\right]}_{<{\rm Tr}\left[{\mathcal{L}}({\mathrm{y}},1,{\bm{V}})^{T}\right]\leq 1}$
    $<{\rm Tr}\mathbb{E}_{{\mathrm{y}}}\left[{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}})^{-1}{\bm{Q}}{\bm{V}}(\mathrm{Cov}[\bm{\mathbf{z}}|y]-{\bm{I}})\right]$
@@
    $\leq\alpha^{-1}\|{\bm{V}}^{\nicefrac{{1}}{{2}}}\|_{\mathrm{op}}^{2}{\rm Tr}({\bm{Q}})=\alpha^{-1}{\rm Tr}({\bm{Q}}),.$
@@
  • Uninformed fixed points. Initializing GAMP with $\bm{M}^{0}={\bm{0}}$, which is a fixed point of the state evolution,
    $\bm{Q}=\alpha\tilde{{\mathcal{F}}}(\bm{Q};,{\bm{V}}),.$ (107)

### v2

@@ C.1 State Evolution
${\mathcal{F}}({\bm{M}};a,{\bm{V}})$ $\coloneqq\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T}(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T})^{-1}{\bm{M}}\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}|{\mathrm{y}}]\right],$ (80)
@@
${\mathcal{G}}({\bm{M}};a,{\bm{V}})$ $\coloneqq\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T}(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T})^{-1}{\bm{M}}\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}|{\mathrm{y}}]{\bm{M}}^{T}(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T})^{-1}{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T}\right].$ (82)
@@
$\langle{\mathcal{F}}({\bm{M}};a,{\bm{V}}),{\bm{M}}^{\prime}\rangle$ $=\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}{\rm Tr}\left[\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}|{\mathrm{y}}]{\bm{M}}^{T}(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T})^{-1}{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T}{\bm{M}}^{\prime}\right]$ (83)
$=\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}{\rm Tr}\left[(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T})^{-1}{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T}{\bm{M}}^{\prime}\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}|{\mathrm{y}}]{\bm{M}}^{T}\right]$ (84)
$={\rm Tr}\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T}(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}^{T})^{-1}{\bm{M}}^{\prime}\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}|{\mathrm{y}}]{\bm{M}}^{T}\right]$ (85)
@@
Furthermore, in order to bound the operator norm of ${\bm{V}}$, we choose
$\gamma_{t}=\|{\bm{V}}^{-1}_{t}+\alpha\mathbb{E}_{{\mathrm{y}}}\left[{\bm{V}}_{t}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}_{t}\left(a-{\bm{V}}_{t}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}_{t}\right)^{-1}\right]\|_{\mathrm{op}},$ (88)
so that $\|{\bm{V}}_{t\to\infty}\|_{\mathrm{op}}=1$.
@@
  • Informed fixed points. (Theorem 3.8) Choosing $a_{t}$ such that
@@
    where we used, introducing the auxiliary notation ${\mathcal{L}}(y,a,{\bm{V}})=((a{\bm{V}}^{-1}-{\bm{V}})\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}|{\mathrm{y}}]+a{\bm{V}}^{-1})^{-1}$
@@
    $=\mathbb{E}_{{\mathrm{y}}}{\rm Tr}\left[{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}})^{-1}{\bm{Q}}{\bm{V}}\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}|{\mathrm{y}}]{\mathcal{L}}({\mathrm{y}},a,{\bm{V}})\right]$
    $\leq\mathbb{E}_{{\mathrm{y}}}{\rm Tr}\left[{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}})^{-1}{\bm{Q}}{\bm{V}}\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}|{\mathrm{y}}]\right]\underbrace{{\rm Tr}\left[{\mathcal{L}}({\mathrm{y}},a,{\bm{V}})\right]}_{<{\rm Tr}\left[{\mathcal{L}}({\mathrm{y}},1,{\bm{I}})\right]\leq 1}$
    $<{\rm Tr}\mathbb{E}_{{\mathrm{y}}}\left[{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}(a-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}})^{-1}{\bm{Q}}{\bm{V}}\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}|{\mathrm{y}}]\right]$
@@
    $\leq\alpha^{-1}\|{\bm{V}}^{\nicefrac{{1}}{{2}}}\|_{\mathrm{op}}^{2}{\rm Tr}({\bm{Q}})=\alpha^{-1}{\rm Tr}({\bm{Q}}).$
@@
  • Uninformed fixed points. (Theorem 3.9) Initializing GAMP with $\bm{M}^{0}={\bm{0}}$, which is a fixed point of the state evolution,
    $\bm{Q}=\alpha\tilde{{\mathcal{F}}}(\bm{Q};a,{\bm{V}}),.$ (92)